OpenStudy (anonymous):

0.7x+6 is equal to or less than 1.5x-5 , the solution set is{x/x ??}

OpenStudy (anonymous):

You have: \[0.7x+6 \le 1.5x-5 \rightarrow 0.7x+11 \le 1.5x\] (add the 5) Subtract the .7x giving: \[11 \le .8x \rightarrow 11 \le \frac{8x}{10}\] Divide by 8/10 or multiply by 10/8 giving: \[\frac{110}{8} \le x\] That simplifies to: \[\frac{55}{4} \le x\] So your solution set would be: \[\left\{ x \in \mathbb{R} |\frac{55}{4} \le x \right\}\]

OpenStudy (anonymous):

As long as my algebra is correct.